OpenStudy (anonymous):

if one angle of a triangle measures 32 degrees and other angle measures 71 degrees, what is the measure of the third angle

OpenStudy (anonymous):

Ok do you know how much a triangle is? in degrees

OpenStudy (anonymous):

in other words the sum of all angles in a triangle should equal what?

OpenStudy (anonymous):

hello are you there?????

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

71+32= 103 180-103= 77 THE THIRD ANGLE IS 77 because the angles can only measure up to 180

OpenStudy (anonymous):

Ok well i will just explain. The sum of all the angles in a triangle should equal 180

OpenStudy (anonymous):

so now we have the equation 71+32+ x=180

OpenStudy (anonymous):

The x is the unknown angle that we need to figure out. Btw the reason i have three numbers is because there are three sides in a triangle

OpenStudy (anonymous):

so 71 +32=103 but it needs to equal 180... So x+103=180..... what plus 103 equals 180?

OpenStudy (anonymous):

Are you there???????

OpenStudy (anonymous):

^ You could've shown her a simpler way, it's not that complicated

OpenStudy (anonymous):

exactly he was making me nervous and sht yelling at me

OpenStudy (anonymous):

I bet, sorry hun

OpenStudy (anonymous):

its ok